### v1

# Training–free AI for Earth Observation Change Detection using Physics Aware Neuromorphic Networks

Abstract
Earth observations from low Earth orbit satellites provide vital information for decision makers to better manage time–sensitive events such as natural disasters. For the data to be most effective for first responders, low latency is required between data capture and its arrival to decision makers. A major bottleneck is in the bandwidth–limited downlinking of the data from satellites to ground stations. One approach to overcome this challenge is to process at least some of the data on-board and prioritise pertinent data to be downlinked. In this work we propose a Physics Aware Neuromorphic Network (PANN) to detect changes caused by natural disasters from a sequence of multi-spectral satellite images and produce a change map, enabling relevant data to be prioritised for downlinking. The PANN used in this study is motivated by physical neural networks comprised of nano-electronic circuit elements known as “memristors” (nonlinear resistors with memory). The weights in the network are dynamic and update in response to varying input signals according to memristor equations of state and electrical circuit conservation laws. The PANN thus generates physics–constrained dynamical output features which are used to detect changes in a natural disaster detection task by applying a distance-based metric. Importantly, this makes the whole model training–free, allowing it to be implemented with minimal computing resources. The PANN was benchmarked against a state-of-the-art AI model and achieved comparable or better results in each natural disaster category. It thus presents a promising solution to the challenge of resource–constrained on-board processing.
keywords:
Neuromorphic AI, Earth Observation, Change Detection
Introduction
Earth Observation (EO) from low Earth orbit satellites provide large volumes of data which are used in a diverse number of applications, including land cover monitoring [1], food security [2] and gas leaks [3]. Machine Learning (ML) techniques have become a critical tool for processing the enormous quantity of EO data [4, 5], with new ML models being designed for change detection tasks that specifically use EO data [6, 7]. An area in which ML using EO data can have a significant impact is the management of natural disasters [8, 9, 10]. For the data to be most effective for first responders, low latency is required between data capture and its arrival to decision makers. One of the largest bottlenecks is the downlinking of data due to the limited bandwidth and coordination of available ground stations [11]. This problem is expected to grow in the future as more data is collected on board satellites due to advances in sensor technology (e.g. hyper-spectral cameras) and an increasing number of satellites [12]. One way to overcome this challenge is to process some of the data on-board and prioritise pertinent data to be downlinked, particularly for time-sensitive applications like natural disaster management. On-board processing will also be critical for future spacecraft deployed to remote sites in the Solar System, or for monitoring solar weather in near-real-time.
Deep ML models, specifically CNNs, have been applied to several natural disaster tasks using EO data, including flood segmentation [13, 14], detecting earthquake affected buildings [15] and volcanic eruptions [16]. To increase the performance of these models, often the number of layers and weights within the model needs to be increased. Chintalapati et al. [17] showed how the architecture of CNN models has evolved to increase performance on an EO image classification task, which resulted in an increase in the number of weights in the models, with the two best performing models, CoCa [18] and BASIC-L [19], using 2.1 and 2.44 billion parameters, respectively. It was noted that most deep learning models have so many parameters that they would exceed the memory budget on most EO satellites. In recent years, work has been done to design smaller AI models specifically for use on-board satellites [20, 21]. Furthermore, missions such as $\Phi$-Sat-1[22] and OPS-SAT [23] have also designed satellites to accommodate AI models on-board. Although these models have been specifically designed for on-board use, all the training still occurs on the ground because of the large computational overhead. Only inference is performed on-board. Since AI models are trained on the ground, they often cannot be trained with the on-board sensor data, leading to degraded performance [24, 25]. Work has been done to retrain models, once operational, with new data as it is acquired [26, 27].
In this study, we introduce the PANN, a AI model with a novel architecture modelled after physical neural networks comprised of nano-electronic circuit elements known as memristors (nonlinear resistors with memory) that mimic synapses [28, 29]. Like biological synapses, the memristive network weights continuously update in response to varying input signals, constrained by physical laws embedded within the model. The PANN, therefore, actively learns as new inputs are fed into the network, such as from a sensor on-board a satellite. As such, the PANN takes a significant shift away from the conventional artificial neural network AI paradigm, as the gradient–descent, back–propagation training of weights is completely replaced by physics–constrained dynamical equations. This is also in contrast to physics–informed neural networks, where the physics is embedded as a component of the model in the artificial neural network paradigm. Typically, this is done by modifying activation functions, gradient descent optimisation techniques, network structure, or loss functions [30, 31]. The added physics normally takes the form of a partial differential equation and is task specific, for example, incompressible Navier-Stokes equations for fluid flow problems [32], whereas with the PANN, the physics is independent of the task.
Physical memristive networks, which the PANN is modelled after, have demonstrated learning tasks such as image classification and sequence memory [33], voice recognition [34, 35], long-/short-term memory and contrastive learning [36], as well as regression tasks of varying complexity [37, 38, 39] and chaotic time series prediction [40]. The PANN model has also been used to demonstrate these and other ML tasks, including transfer learning and multi-task learning [41, 42, 43, 44, 45, 46, 47, 48]. These ML tasks were implemented using reservoir computing, where only weights in a single output layer need to be trained. This approach offers a substantial advantage over deep neural network models as it does not need computationally intensive training algorithms and thus does not consume as much energy. These properties make the PANN an ideal candidate for use on-board satellites.
In this study, we use a natural disaster dataset released by Růžička et al. [21] to evaluate the performance of the PANN model on a change detection problem, that contains a time series of multi-spectral EO images from Sentinel-2. We benchmark the PANN against a state-of-the-art deep learning model and achieve comparable or better results. Additionally, we use a distance measurement to detect changes, bypassing the need for an output layer altogether. Hence, we do not need to train any part of the model and the PANN does not need to be trained on the ground prior to deployment.
This article is organised as follows: in “Results” we show the results from the PANN model and compare them to a deep learning model for the same change detection task. We also show visualisations of the features the PANN model extracts. In “Discussion” we outline key findings from the results and their implication for use on-board satellites. Finally, “Methods” presents the details of the PANN model and workflow.
Results
The PANN model is used here to detect regions of change from natural disasters using EO multi-spectral images. The change detection task requires the model to take in a sequence of five images, where the final image is after a natural disaster event. The model outputs a change feature map that highlights the regions affected by the natural disaster, by breaking the images into smaller tiles and assigning a change likelihood score for each tile. The PANN model used in this study is an adaptive network, with weights that adjust with each new input. This allows the PANN to generate features from the inputs without requiring any training (see Methods for details). To determine if an image area has changed, the distance between inputs are compared in the feature space rather than the input space (pixel comparison). By applying a distance-based metric directly to the feature space, no training is performed at any stage in the model. The dataset includes four different categories of natural disaster events: fires, floods, hurricanes and landslides. Additionally, each natural disaster event is accompanied by a change mask which is used only to evaluate the performance of the PANN model.
Figure 1 shows the output change maps produced by the PANN model for each type of natural disaster event. Alongside the change maps are the RGB images directly before and after the natural disaster events and the associated change mask, which also includes clouds from the images directly before and after the natural disaster event. Although only one image is shown before the natural disaster event in Fig. 1, a total of four images before the natural disaster event were used to create the displayed change map. For the four natural disaster events, the main areas of change have been correctly identified as changes in the change maps, with a limited amount of noise in the images being considered as a change, such as the vegetation colour change in the fire natural disaster event (top row), or the sediment being washed into the ocean in the flood natural disaster event (second row). Clouds are detected as changes, but are masked out of the change maps shown in Fig. 1, since they are ignored when evaluating the change maps to allow for a more direct comparison to a state-of-the-art deep learning model (see Methods for details).
Figure 1: RGB images (columns 1 and 2) directly before and after a natural disaster event, the corresponding target change mask (column 3) and the output change map produced by the PANN model (column 4) for each type of natural disaster: Top row – an example fire natural disaster event; Second row – an example flood natural disaster event; Third row – an example hurricane natural disaster event; Bottom row – an example landslide natural disaster event. Cloud-cover has been masked from the images and corresponds to yellow in the change mask, with genuine differences coloured aqua.
For each natural disaster event a change map was created, with the order in which the natural disasters were passed into the PANN model being randomised. We evaluated the quality of the change maps by calculating the Area Under the Precision—Recall Curve (AUPRC). The change likelihood score for each tile is given by equation (3) in Methods, which depends on the distance metric used to measure the position of the tiles in the feature space. Table 1 reports the overall AUPRC as a percentage for each of the natural disaster categories using three different distance metrics, Euclidean, Cosine and Correlation distances. The Correlation distance significantly outperforms the other distance metrics and is also much more consistent across the different natural disaster classes.
Distance Metric Fires Floods Hurricanes Landslides
Euclidean $90.084\pm 0.007$ $42.79\pm 0.04$ $43.1\pm 0.07$ $27.38\pm 0.09$
Cosine $86.71\pm 0.01$ $39.36\pm 0.03$ $22.50\pm 0.01$ $35.2\pm 0.2$
Correlation $\mathbf{93.66\pm 0.01}$ $\mathbf{59.26\pm 0.02}$ $\mathbf{76.51\pm 0.05}$ $\mathbf{73.5\pm 0.2}$
Table 1: Comparison between different distance metrics used to calculate the PANN change score for different natural disaster classes. All values are calculated as mean AUPRC (%) $\pm$±SEM over 5 runs. Best scores are highlighted.
We benchmark our PANN model against a state-of-the-art variational autoencoder model [49], called RaVAEn, which was specifically designed to be a relatively small AI model that could be deployed on-board satellites. We also compare the PANN model to a baseline method that uses the same change score equation but compares the different tiles in the pixel space rather than the feature space like the other two AI models. When comparing models, it should be noted that the PANN has an automatic feature engineering process incorporated into the model setup that is not present in the RaVAEn model or the baseline method. Table 2 reports the performance of the three different models for each of the natural disaster classes. Both AI models, RaVAEn and PANN, outperform the baseline method in all classes. The RaVAEn and PANN models achieve comparable results in the fires, hurricanes and landslides, with the PANN slightly ahead for the fires and hurricanes, while RaVAEn is slightly ahead for the landslides. With the flood natural disasters, the PANN achieves the highest results by a significant margin with a score 14% higher than that of RaVAEn.
Model Fires Floods Hurricanes Landslides
Baseline 86.5 37.6 58.0 62.9
RaVAEn $90.98\pm 0.08$ $44.5\pm 0.7$ $76.10\pm 0.08$ $\mathbf{76.5\pm 0.8}$
PANN $\mathbf{93.66\pm 0.01}$ $\mathbf{59.26\pm 0.02}$ $\mathbf{76.51\pm 0.05}$ $73.5\pm 0.2$
Table 2: Comparison between change scores predicted by different models for different natural disaster classes. All values are calculated as AUPRC (%), with values for the PANN model corresponding to the best scores from Table 1.
Feature Space
The AI models RaVAEn and PANN outperform the baseline method, since they can extract the most relevant features from the tiles and group similar tiles together, while keeping dissimilar tiles further apart in their feature spaces. To visualise the high dimensional feature space of the PANN, we used the Uniform Manifold Approximation and Projection (UMAP) [50] method to reduce the dimensions of the feature space into 2D. The UMAP used 20 neighbours, a minimum distance of 0.4 and the correlation distance, unless otherwise stated. Figure 2 shows the UMAP 2D projection of the feature space for a flooding event. The left-hand side shows the position of the tiles in the image directly before and after the flooding event. For tiles after the flooding event, the percentage change in each tile is denoted by the colour bar. The right-hand side of the figure shows the same projection of the feature space but with the tile images instead of points. The change tiles are grouped together at the bottom of the projections and the remaining tiles are also sorted according to the features in the tiles, with greener farmland tiles on the left-hand side and tiles with the white objects on the right. Tiles that have missing values, from the north alignment of the swath, are also grouped towards the bottom right, with the tiles that have the missing values across the top being separated from the tiles with the missing values on the right-side of the tile. For comparison with the UMAP projection for the pixel space, see Supplementary Fig. 8.
Figure 2: 2D UMAP projections of the PANN feature space, using the correlation distance for a flood natural disaster event: Left – positions of before (blue) and after (orange) tiles, with percentage change in the after tiles indicated by the colour bar; Right – the corresponding tile images. The tiles with flooding occupy a distinct arc-like cluster at the bottom of the projection.
Like most dimensionality reduction techniques, UMAP aims to reduce the number of dimensions while maintaining the overall structure and information of the original, higher dimensional space and thus depends on the distance metric used to measure the distances between points. Figure 3 shows the UMAP projection of the feature space for a hurricane natural disaster for the three distance metrics used in Table 1 and shows how the clustering of the tiles varies depending on the distance metric being used. This provides a visual qualitative tool to understand the large range in results reported in Table 1 due to the change in distance metric being used.
Figure 3: Visual comparison between the 2D UMAP projection of the PANN feature space for a hurricane natural disaster using three different distance metrics: Euclidean, Cosine and Correlation distances.
Discussion
In this study, we introduced the PANN model for a natural disaster change detection task and compared it against a state-of-the-art AI model RaVAEn and a baseline method. All three models receive the same multi-spectral EO images for the inputs and use the same change detection equation given by equation (3), for each tile in the image. The difference in the model performances, therefore, is attributed to their ability to extract the most relevant features from the inputs, while ignoring noise within the inputs. We note, however, that while the PANN model receives the same inputs as the other models, only the inputs corresponding to the bands selected by the automatic feature engineering process, on a scene by scene basis, are passed to the relevant individual networks. This inherently allows the PANN model to ignore irrelevant information as determined by the domain knowledge embedded in the feature engineering process. The baseline method does not extract any features from the inputs and instead does the comparison directly in the pixel space. This makes the baseline method more sensitive to noise in the images and, as expected, it exhibits the lowest performance as reported in Table 2. The machine learning models, on the other hand, extract meaningful features from the tiles and hence outperform the baseline method in every natural disaster category. The PANN and RaVAEn models achieve a comparable result in three of four of the natural disaster categories, demonstrating that both models extract equally useful features for these categories. For flood natural disasters, however, the PANN model outperforms the RaVAEn model and thus arguably extracts more meaningful features.
While being able to extract the features is the most important task for the PANN model, how the feature space is navigated is also important for the performance on the PANN model. This is seen in Table 1 where the same feature space can produce significantly different results based on the distance metric used to navigate the feature space. This is visualised in the UMAP projections in Fig. 3, where different distance metrics are used to create the projections. The feature space produced by the PANN has a high dimensionality, as such the relative Euclidean distance between points tends to decrease[51]. This would result in the change likelihood score decreasing, particularly for change tiles and is likely why the Euclidean distance does not produce the best results. What is surprising is the large difference between the Cosine and Correlation distances, given that both metrics are angle-based measures. The Cosine distance is given by equation (4) and measures the angle between the vectors in the feature space. The Correlation distance is the same as the Cosine distance except each sample is mean centred first, as given by equation (5). The mean-correction for the Correlation distance effectively changes the location of the origin and hence the angle between the vectors also changes [52]. One possible reason why the mean correction has such a large impact on the performance is that it helps reduce covariate shift in the output sequence that can arise from the dynamical nature of the PANN. The covariate shift effect can be seen in Fig. 3 (and Supplementary Fig. 9), where there is a much greater overlap of points at the different timesteps (that are not changes) in the Correlation projection compared to that of the Cosine projection. Interestingly, batch normalisation layers were introduced to deep neural networks to help reduce the covariate shift between layers in the networks [53] and have been widely accepted to speed up training times and increase performance. We note that the PANN does not have any batch normalisation layers, while the RaVAEn model does, and that the best performing metrics for the models are the Correlation and Cosine distance, respectively.
Our results show that the PANN model achieves comparable-to-better results than other state-of-the-art AI models for the natural disaster change detection task conducted in this study. The more impressive result here is that the PANN was able to achieve this result without any training to create the mapping from the input space to the feature space. The PANN is training-free as it continuously adapts with each new input. The model equations are relatively light–weight and thus, evaluation of the PANN model was able to be carried out using only CPUs in this study. The physics–based equations that constitute the PANN model are a significant deviation from conventional deep multi-layered bipartite neural networks that typically use mathematical activation functions to perform nonlinear transformations and stochastic gradient decent with backpropagation as the learning algorithm to train networks weights into an optimal static state. As such, the iterative training process is computationally expensive and usually requires cloud-based GPUs [54], to produce a static model.
The training-free nature of PANN makes it an ideal candidate for extreme edge computing applications, particularly for EO analysis on-board satellites. Specifically, there are two main challenges that are inherently addressed by the training-free nature of the PANN. The first is that ML models typically do not generalise well and need to be trained for a specific sensor. When ML models are trained using data from a different sensor, there is often a notable decrease in performance [24, 25]. Furthermore, training a ML model with only simulation data can be difficult [17]. The training-free nature of the PANN therefore makes it a promising candidate to circumvent this issue. The second is that ML models are often retrained as new data from the desired sensor become available to increase performance and to account for any data-shift problems. This process requires the new data to be downlinked, to retrain the model on the ground before uplinking the updated model to the satellite [26]. The dynamical nature of the PANN would allow this whole process to be bypassed as the PANN inherently updates and learns from each new input from the sensor.
In conclusion, we demonstrated a training-free AI model, the PANN, for a natural disaster EO change detection task that contained four types of natural disasters: fires, floods, hurricanes and landslides. We benchmarked our model against a state-of-the-art AI model and achieved better results for floods and comparable results in the remaining three categories, without any training. We found that the PANN was able to achieve these results due to its ability to extract meaningful features from the inputs and sort them accordingly in the feature space. Being able to navigate the feature space, therefore, plays a vital role in the model’s performance. The training-free nature of the PANN makes it an ideal candidate for use on-board satellites, processing EO data at the extreme edge.
Methods
Data
The images are level 1C pre–processed multi-spectral images from the Multi-Spectral Imager (MSI) camera on-board the Sentinel-2 Earth observation satellite constellation. Only the 10 highest spatial resolution bands are used in this study. Table 3 shows the specific bands used in this study along with the corresponding index number used to reference the bands. The dataset covers four categories of natural disasters: fires, floods, hurricanes and landslides. Each category includes five natural disaster events, except for floods which have four, giving a total of 19 natural disaster events. For each natural disaster event there are five sequential images; the first four are all before the natural disaster event and the final image is after. Along with each natural disaster event, there is an accompanying change mask that contains three labels: unaffected regions, affected areas, and clouds from the fourth and fifth images. An example of a hurricane disaster event is shown in Fig. 4, where damage to the island vegetation can be seen in the fifth image and the change mask image shows the unaffected regions in purple, the affected region in aqua and the clouds in yellow. The change mask in this study is only used to evaluate the performance of the proposed model.
Band B2 B3 B4 B5 B6 B7 B8 B8a B11 B12
Index (m) 01 02 03 04 05 06 07 08 09 10
Table 3: Sentinel-2 bands used in this study and the corresponding index used to reference each band.
The data was preprocesessed by first taking the log of the rgb pixel values and then scaling the values to the interval [-1, +1] for each band in the data. This process is described by equation (1):
$x^{\dagger}=\log(x)\qquad,\qquad x^{\ddagger}=2\quantity(\frac{x^{\dagger}-\mbox{min}(x^{\dagger})}{\mbox{max}(x^{\dagger})-\mbox{min}(x^{\dagger})})-1$ (1)
where the min and max values for each band were chosen to be consistent with Ref. [21], to allow a direct comparison with their state-of-the-art variational autoencoder model. Missing values were set to a value just above zero of 0.005.
Figure 4: An example of a hurricane natural disaster scene, with the first 4 images before the hurricane and image 5 is after. The change mask (last column) shows the affected (aqua) and unaffected (purple) regions along with the clouds (yellow) in the images directly before and after the hurricane.
Model Architecture
The PANN is based on a physically–motivated model of a physical neuromorphic network comprised of self–organised nanowire networks [41, 43, 45]. This is done by first creating a nanowire network and then converting it into a graph representation which is used in the PANN. This process is shown in Fig. 5. In this study, the network for the PANN was created by distributing 803 nanowires over a $158\times 158\,\mu$m 2D plane with nanowire centres sampled from a generalised normal distribution with a beta value of 5 and with nanowire orientations sampled on $[0,\pi]$. Nanowire lengths were sampled from a Gaussian distribution with an average of $30\,\mu$m and a standard deviation of $6\,\mu$m. As nanowire networks are nano-electronic systems, input signals are delivered via electrodes, which were modelled with a diameter $5\,\mu$m and with evenly spaced placements in a $16\times 16$ grid over the 2D plane with a margin of $15\,\mu$m, giving a pitch of $8\,\mu$m between electrodes. Where the nanowires overlap with other nanowires or electrodes, electrical junctions are formed whose internal states evolve in time, as described below. Figure 5(a) shows the physical nanowire network, with the nanowires displayed as grey lines and nanowire–nanowire junctions as small grey dots. The electrodes are shown as the large green circles and the nanowire–electrode junctions are shown as small dark green dots. Figure 5(b) shows the corresponding graphical representation of the network used in the PANN. The black nodes correspond to the nanowires and the green nodes correspond to the electrodes, used as input nodes. The edges correspond to the electrical junctions. The resulting network had a total of 1059 nodes and 12,279 edges. 256 nodes were selected as input nodes and from the remaining 803 nodes, 400 were selected at random as output nodes (note that in a real physical device, electrodes would also serve as output nodes).
Figure 5: (a) PANN model nanowire representation. The nanowires are the grey lines and the electrodes are the large green circles. Nanowire–nanowire junctions are shown as grey dots and nanowire–electrode junctions are shown as dark green dots. (b) graph representation of the PANN. Green nodes are the input nodes which correspond to the electrodes. Black nodes correspond to the nanowires and edges correspond to the junctions, nanowire–nanowire and nanowire–electrode junctions.
The network in the PANN has a heterogeneous, neuromorphic topology and operates like a complex electrical circuit with nonlinear circuit components known as memristors [55, 45]. As such, inputs are treated as input voltage signals and Kirchhoff’s conservation laws are solved at each time step to calculate the node voltage distribution across the network. The edge weights are conductances that evolve in time according to a memristor equation of state:
[…]
with all edges set to an initial state $\lambda(t=0)=0$. In this study, the following parameters were used: $V_{\text{reset}}=5\times 10^{-3}\,$, $V_{\text{set}}=1\times 10^{-2}\,$ and $\lambda_{\text{max}}=1.5\times 10^{-2}\,$ s. Full details of the model and its validation against experimental nanowire networks are provided in Refs. [43, 45]. In contrast to artificial neural networks, the dynamic nature of these physical neuromorphic networks means that weights are not trained, which not only makes them more energy efficient, but also allows all the data to be used for evaluating the model.
Model Setup
Figure 6 shows the process of feeding the natural disaster image sequence into the PANN model. First an automatic feature engineering step determines which bands to use for that specific natural disaster event. This is motivated by a physical understanding of the event: for example, water exhibits the greatest contrast compared to land at infrared wavelengths. The images are then subdivided into a series of $32\times 32$ pixel sequence tiles, $x^{a,b}_{N}(t)$, where $a,b$ is the location of the tile and $N\leq 10$ is the number of bands as determined by the feature engineering process. The images are tiled column by column starting at the top left hand corner. The tile sequence is passed through a max pooling layer that has a pooling size (2,2), a stride of 2 and does not use any padding, hence halving the spatial resolution of the tile sequence to $16\times 16$ pixels. The tile sequence is then split into individual bands, creating the input signals $U^{a,b}_{m}(t)$, where $m$ is the index to each PANN for each spectral band as given in Table 3. As such, there are a total of 10 identical PANN networks, each taking inputs for a specific band. The readout values of the PANNs used for the given tile sequence are concatenated to give the final features, $F^{a,b}(t)$ for that tile location. Hence, not all 10 PANNs are used for every natural disaster event, as demonstrated in Fig. 6, where only PANN numbers 01-03 and 09 are used and the greyed out PANNs (PANN 04 – 08 and 10) are not used for this natural disaster event.
Figure 6: Diagram of the PANN workflow showing the pipeline from the input images to the output features, $F^{a,b}(t)$ from the networks. First, automatic feature engineering is applied to the input images to determine which bands of the images to use. The images are then broken into a sequence of small $32\times 32$ pixel tiles, $x^{a,b}_{N}(t)$, where $a,b$ is the location of the tile and $N$ is the number of bands. The creation of the tile sequence is shown by the coloured (red, orange, yellow, lime and green) boxes and are then fed into the max pooling layer. The bands are separated, as denoted by $\bigotimes$, with each band being fed into a unique PANN model. In this example PANN01-03 and PANN09 are used. The outputs from all the PANNs that are used are concatenated together, as denoted by $\bigoplus$, to produce the feature sequence $F^{a,b}(t)$ for tile $x^{a,b}_{N}(t)$.
To determine whether a change from a natural disaster occurred at a given tile location, the distance between the images of the tile sequence is measured in the feature space. It is assumed that if there is any change it is in the last image of the sequence, so only the distances between the final image and each of the previous images is compared. This is the same change score method as introduced in Ref. [21]. Formally, the change score, $S$, is given by
$S(x^{a,b}(t))=\min_{i=1,2,3,4}\left[\mbox{dist}(F^{a,b}(t-i),F^{a,b}(t))\right]\qquad,$ (3)
where $dist$ is an arbitrary distance metric. In this study we focus on three different distance metrics: Euclidean, Cosine and Correlation. Unless otherwise stated, the Correlation distance was used. The Cosine distance is an angle-base metric and is given by
$\text{CosineDist}(\mathbf{u},\mathbf{v})=1-\frac{\mathbf{u}\cdot\mathbf{v}}{||\mathbf{u}||*||\mathbf{v}||}\qquad.$ (4)
The Correlation distance is the mean corrected version of the Cosine distance, given by
$\text{CorrelationDist}(\mathbf{u},\mathbf{v})=\text{CosineDist}(\mathbf{u}-\text{mean}(\mathbf{u}),\mathbf{v}-\text{mean}(\mathbf{v}))\qquad.$ (5)
Distances between the tile sequence images are compared directly in the features space, so that it is not necessary to learn a mapping function from the feature space to a target variable. This means it is not necessary to train any part of the model.
To evaluate the quality of the change maps produced by the PANN we use the AUPRC [56]. To calculate the curve, we treated each pixel in the output change maps as a positive or negative example, across all natural disaster events in that category. As such, a separate precision-recall curve was produced for each class of natural disaster and an AUPRC value. Areas labelled as clouds in the change mask were ignored when calculating the AUPRC. This was done to be consistent with Ref. [21] and to allow for a more accurate comparison between the PANN and RaVAEn models.
Automatic Feature Engineering
The automatic feature engineering process is performed once for each natural disaster scene and is performed first, before the normalisation and tiling stages. The feature engineering process leverages common indices such as the normalised burn ratio index [57] and the normalised difference vegetation index [58] to calculate natural disaster class scores. Using these scores, a binary decision tree is followed, as shown in Fig. 7, to determine what class of natural disaster is in the scene. The score for each natural disaster class is calculated by creating an index image for the images directly before and after the natural disaster event using equation (6), where $BX$ and $BY$ are predetermined bands depending on what type of natural disaster is being assessed. The images are converted to a binary image using a predefined threshold value and a difference image is created by a pixel–wise comparison. The final score is the percentage of pixels that record a change. Once a natural disaster type has been selected, the feature engineering process then selects a subset of bands to be used in the rest of the model depending on the natural disaster class. The specific values and bands used for the natural disaster class score are shown in Fig. 7.
$\frac{BX-BY}{BX+BY}$ (6)
Figure 7: Decision tree for the automatic feature engineering step. At each branch a natural disaster score is calculated for the different categories. For each score the bands and threshold values used to create the index image are given. The bands used for each natural disaster type once selected are also shown.
We note that more sophisticated index images [59, 60] could be used to categorise the natural disasters and potentially further improve results. Furthermore, similar feature engineering processes could be used to tune the model to detect changes from specific natural disasters or even clouds. The focus here is to show that the simplest automatic feature engineering process, one that does not require training, can be used to easily incorporate domain knowledge and improve the performance of the model.
Data Availability
We are releasing the full code alongside this paper at https://github.com/ssmi9157/ChangeDetectionPANN. The dataset analysed in this study is available in the RaVAEn repository, https://github.com/spaceml-org/RaVAEn.
Acknowledgements
The authors acknowledge the use of the Artemis High Performance Computing resource at the Sydney Informatics Hub, a Core Research Facility of the University of Sydney. S.S. is supported by an Australian Government Research Training Program (RTP) Scholarship.
Author contributions statement
S.S. and Z.K. designed the experiments. S.S. performed the experiments with guidance from C.P. and Z.K. S.S., C.P. and Z.K. analysed the results. S.S. drafted the manuscript with consultation from Z.K. All authors critically reviewed the manuscript. Z.K. supervised the project.
Additional information
Competing interests
Z.K. is a founder and equity holder of Emergentia, Inc., which filed a non-provisional U.S. patent application (no. 18/334,243) for the software simulator described in this work. C.P. was employed by Trillium Technologies during this work. The other author declares no competing interests.
Appendix
Figure 8: The 2D UMAP projection of the pixel space for a flood natural disaster event: Left – position of the before (blue) and after (orange) tiles and with percentage change in the after tiles indicated by the colour bar; Right – the corresponding tile images.
Figure 9: Visual comparison between the 2D UMAP projection of the PANN feature space for a hurricane natural disaster, for the full tile sequence, using three different distance metrics: Euclidean, Cosine and Correlation distances.